OpenStudy (anonymous):

MEDAL! ~PROBABLITY~ Alice and Bill are planning to have three children. (Assume it is equally likely for a boy or a girl to be born.) What is the probability that all three will be the same gender?

OpenStudy (toasty_the_robot):

1/3

OpenStudy (anonymous):

1/3

OpenStudy (toasty_the_robot):

well technically 1.5/3

OpenStudy (anonymous):

can someone explain to me why it would not be 1/8?

OpenStudy (mathmate):

For one child, probability is 1/2 for a boy. For two boys, probability is (1/2)^2=1/4. For three boys, probability is (1/2)^3=1/8 = P(BBB) Assuming equality of probability between boys and girls, probability of three girls is also = 1/8 = P(GGG) So probability of having 3 children of same gender = P(BBB)+P(GGG)

OpenStudy (anonymous):

so was i correct at 1/8? or no? hahaha

OpenStudy (mathmate):

Calculate P(BBB)+P(GGG) and you will find out.